solve for the proportion of the area of the following probability notations then sketch the normal curve shading the required region of each given notation




1. P (-2 < z <1)



2. P (z > 3)



3. P (z < -2)



4. P (z > -1.5)



5. P (z = -1)




1
Expert's answer
2022-02-24T08:44:03-0500

1.

"P (-2 < z <1)=P(z<1)-P(z\le -2)"

"\approx0.84134-0.02275\approx0.8186"

2.

"P (z>3)=1-P(z\le3)"

"\approx0.00135"



3.

"P ( z <-2)\approx0.02275"

4.

"P (z>-1.5)=1-P(z\le-1.5)"

"\approx0.9332"



5.

"P (z=-1)=0"